Suppose x = a and x = e are two solutions to the quadratic equations x2 + px = q. Use Vietes method to find the relationshipSuppose x = a, e are two solutions to the cubic equation x3 + px = q . Use Vietes approach to find the relationship between= Suppose x = a is a solution to x3 +p=qx , while x = e is a solution to x3 approach to find the relationship between a, e. q
Answer #1

Ques x²+ px =q x² + p x-q=0 x = a je reools of ove this equation. By Nietes method , ate =- There are 2 eloctions. ae =-q K.from 6 & ② - Cate) = late) q+aelate) = 0 H* These are the 3 relations. Que 3 x² = qx + ? (2-93-p2 het (2=ki be the red 3 root